# Joint semi-supervised and contrastive learning enables zero-shot domain-adaptation and multi-domain segmentation

### v1

@@ Abstract
Despite their effectiveness, current deep learning models face challenges with images coming from different domains with varying appearance and content. We introduce SegCLR, a versatile framework designed to segment volumetric images across different domains, employing supervised and contrastive learning simultaneously to effectively learn from both labeled and unlabeled data. We demonstrate the superior performance of SegCLR through a comprehensive evaluation involving three diverse clinical datasets of retinal fluid segmentation in 3D Optical Coherence Tomography (OCT), various network configurations, and verification across 10 different network initializations. In an unsupervised domain adaptation context, SegCLR achieves results on par with a supervised upper-bound model trained on the intended target domain. Notably, we discover that the segmentation performance of SegCLR framework is marginally impacted by the abundance of unlabeled data from the target domain, thereby we also propose an effective zero-shot domain adaptation extension of SegCLR, eliminating the need for any target domain information. This shows that our proposed addition of contrastive loss in standard supervised training for segmentation leads to superior models, inherently more generalizable to both in- and out-of-domain test data. We additionally propose a pragmatic solution for SegCLR deployment in realistic scenarios with multiple domains containing labeled data. Accordingly, our framework pushes the boundaries of deep-learning based segmentation in multi-domain applications, regardless of data availability —– labeled, unlabeled, or nonexistent.
@@ I Introduction
Domain adaptation techniques aim to maintain a similar performance on images from new domains while reducing the annotation effort required in such newer datasets. Unsupervised domain adaptation aims to leverage information learned from a labeled data domain for applications in other domains where only unlabeled data is available. To this end, many deep learning methods have been proposed [8], mostly using generative adversarial networks, e.g., to translate visual appearance across OCT devices [9]. Ultimately, it is desirable for learned models to implicitly generalize to other domains without requiring any new data, i.e., ideally not even unlabeled data from a new domain. This scenario, known as zero-shot domain adaptation [10], provides huge advantages in real-world deployment by obviating any retraining or customization being required per test domain encountered.
Advances in self-supervised learning have been increasingly successful in extracting informative features from images without accessing their labels, hence enabling the exploitation of much larger unlabeled datasets [11, 12, 13, 14, 15, 16, 17, 18]. These methods usually employ different versions of Siamese networks [19] that can be applied to different inputs in parallel while sharing their weights. For self-supervision, typically different augmentations of an image are treated as positive samples, which are desired to map close-by or to a single point on a learned manifold, and therefore minimizing the feature distances between these positive samples when passed through a Siamese network. To prevent such optimization from collapsing to a constant output, different approaches have been proposed: For instance, SimSiam [14] employs a stop-gradient operation for one of the Siamese encoders. Another successful strategy, contrastive learning (CL), introduces the concept of negative pairs, the distances of which should meanwhile be maximized. SimCLR [12] is one of the most widely-adopted CL frameworks. Self-supervised learning is commonly used for model pretraining, typically based on natural images such as ImageNet [20]. Such pretrained model is subsequently fine-tuned or distilled for downstream tasks such as classification, detection, or segmentation [13].
@@
We propose SegCLR as a CL solution to improve segmentation quality by leveraging both manual annotations and unlabeled data. To that end we introduce a semi-supervised framework for joint training of CL together with segmentation labels. Preliminary results of this work have been presented as a conference contribution [23] for two datasets acquired with different OCT scanners, where the domain shift originates from a change in acquisition device. SegCLR without any target-domain labels is shown to achieve unsupervised domain adaptation results close to an upper bound that uses supervised learning on the target domain. Our additional contributions in this work are: (1) In addition to acquisition device change, we present results with a third dataset with a new disease indication, which helps demonstrate the benefits of SegCLR also for domain shift in disease type. (2) An analysis of the amount of unlabeled target data is presented, indicating its limited effect on the resulting segmentation performance. (3) Exploiting the finding above, we propose a zero-shot domain adaptation use-case of our SegCLR framework, for which an extensive analysis reveals substantial benefits showing that SegCLR trained models are implicitly better at generalizing to unseen domains. (4) An analysis of the stability of the results upon distinct network initialisations reveals that most proposed SegCLR variants perform well and similarly, confirming the stability of the overall SegCLR framework to method setting choices such as pair generation and contrastive projection strategies. This further highlights the importance of analyzing many randomized result replicates in assessing deep learning models in general. (5) In a multi-domain training scenario of SegCLR, we show that with labeled data available from multiple domains, SegCLR applied to the domains simultaneously yields results superior to conventional supervised training as well as SegCLR trained on individual domains, potentially better leveraging cross-domain information.
@@ II Methods
The considered learning frameworks are illustrated in Fig. 1. Let a source domain ($D^{s}$) contain labeled data, with which Baseline models can be trained following typical supervised learning schemes. Distinct domains for which only unlabeled images can be observed are considered target domains ($D^{t}$). We first study an unsupervised domain adaptation setting as in [23], where SegCLR is trained for a source domain $D^{s}$ and a target domain $D^{t}$ by using a combination of supervision and contrastive losses. We then investigate its zero-shot domain adaptation application, where SegCLR is trained using only data from a labeled source domain $D^{s}$, while being applied on unseen target domains $D^{t}$, i.e., on domains for which no data (labeled or unlabeled) is used in training.
Figure 1: Illustration of Unsupervised Domain Adaptation (UDA) and Zero-Shot Domain Adaptation frameworks studied herein for SegCLR. The colored ellipses indicate the losses to use in training, with the variable superscripts representing the domain being source (s) or target (t).
@@ II-A Supervised learning
For the segmentation backbone, we adopted the proven UNet architecture [2] (details in Fig. 2), which can be modeled as $F(\cdot)$ processing an image $x$ to predict a segmentation map $p=F(x)$ to approximate an (expert-annotated) ground truth segmentation $y$. $F$ is learned by minimizing a supervised loss $\mathcal{L}_{\mathrm{sup}}$, which in our work is the logarithmic Dice loss of labeled data in a source domain $D^{s}$:
$\mathcal{L}_{\mathrm{sup}}=-\sum_{(p_{i},y_{i})\in D^{s}}\log{\frac{2\sum_{j\in\mathrm{pixels}}y_{i}^{j}p_{i}^{j}}{\epsilon+\sum_{j\in\mathrm{pixels}}(y_{i}^{j}+p_{i}^{j})}}$ (1)
for all training images $(x_{i},y_{i})\in D^{s}$, where $\epsilon$ is a small number to avoid division by 0.
@@ II-B Self-supervised learning
Self-supervised learning aims to optimize an encoder $E(\cdot)$ to achieve good representations $h=E(x)$ without the need of manually annotated labels $y$. In the literature (for mostly classification tasks), the ResNet architecture is commonly used for $E(\cdot)$. Instead, we use the existing UNet encoder in order to adapt the learned features $h$ for our intended segmentation task. A subsequent contrastive projection head $C(\cdot)$ then maps the bottleneck-layer features to vector projections $z=C(h)$ on which the contrastive loss $\mathcal{L}_{\mathrm{con}}$ is applied. The architecture used for $E(\cdot)$ and $C(\cdot)$ is illustrated in Fig. 2. We herein employ two widely adopted self-supervised learning frameworks: SimCLR [12] and SimSiam [14].
@@ II-C Joint semi-supervised and contrastive learning
We adapt the USCL joint training strategy, which was proposed for US video classification, to our segmentation task on 3D images by combining $\mathcal{L}_{\mathrm{sup}}$ and $\mathcal{L}_{\mathrm{con}}$ in a semi-supervised framework illustrated in Fig. 3. Considering a source domain $D^{\mathrm{s}}$ and a target domain $D^{\mathrm{t}}$, total loss $\mathcal{L}$ is calculated as follows:
@@
Figure 3: Illustration of semi-supervised contrastive learning framework for unsupervised domain adaptation. The SegCLR block corresponds to the architecture in Fig. 2. The repel losses are not used by SimSiam. Supervised losses are only used when labeled images exist. While this framework is flexible to accommodate any number of labeled images, at least one is required to drive the decoder arm of the underlying UNet.
@@ II-D Pair generation strategies
The approach chosen for contrastive pair generation is key in self-supervised learning. We herein propose and compare different pair generation functions $P(\cdot)$ for volumetric OCT images, as illustrated in Fig. 4.
First, the augmentation-based pair formation typically employed for natural images (e.g., in SimCLR and SimSiam) is adapted for volumetric OCT images, denoted as $P_{\mathrm{a}}$. To that end, labeled slices in $D^{\mathrm{s}}$ and random slices in $D^{\mathrm{t}}$ are augmented with horizontal flipping ($p=0.5$), horizontal and vertical translation (within 25% of the image size), zoom in (up to 50%), and color distortion (brightness up to 60% and jittering up to 20%). For color augmentation, images are first transformed to RGB and then back to grayscale.
@@ II-F Implementation
Adam optimizer [24] was used in all models, with a learning rate of $10^{-3}$. Dropout with $p=0.5$ is applied on the layers illustrated with orange arrows in Fig. 2. The projection head $\rho^{\textrm{MLP}}$ in $C(\cdot)$ is formed by two fully-connected layer with 128 units each, where the first one uses group normalization [25] (with a group size of 4) and ReLU activation. We heuristically set $\lambda=20$ and the standard deviation for $P_{\mathrm{s}}$ as $\sigma=250\,\mu m$, which is the range for which we observe roughly similar features across slices. Supervised models were trained for 200 epochs, and the model at the epoch with the highest average Dice coefficient across classes on the validation set was selected for evaluation on a holdout test set. Our implementation is in Tensorflow 2.7, ran on an NVIDIA V100 GPU.
@@ II-G Evaluation setup
[…]
Averaging metrics across classes with a large variation may lead to bias. Thus, for the metrics reported as boxplots, each per-slice metric $m^{c}_{i}$ for method $i$ and class $c$ is first normalized by its class Baseline $m^{c}_{\textrm{bas}}$, and then these are averaged over all $c$ and images on the test set. The average of all Baseline replicates is used as $m^{c}_{\textrm{bas}}$ for calculating the relative metrics of all other models. Absolute metrics for each class are also reported in the tables as mean $\pm$ standard deviation. To facilitate model comparison, the metrics included in most tables in the main text are averaged across classes. To further allow for the interpretation and comparison of absolute segmentation performances, we included detailed tables in Supplementary Materials with metrics separately for individual labels.
For comparing and ranking models, we adapted the strategy from [26] to simultaneously consider both Dice and UVD as complementary metrics. Accordingly, the models are first ranked among each other on individual B-scans, separately based on their Dice and UVD. These rankings are next averaged across the B-scans and the two metrics, to obtain and report a final ranking of each model for each domain.
@@ III Dataset
All slices (B-scans) from the three datasets were resampled to $512\times 512$ pixels with the approximate resolution of $10\times 4$ $\mu$m/pixel. Selected B-scans from $D_{1}$ and $D_{2}$ were manually annotated for fluid regions of potential diagnostic value, i.e., intraretinal fluid (IRF), subretinal fluid (SRF), pigment epithelial detachment (PED), and subretinal hyperreflective material (SHRM). B-scans from $D_{3}$ were annotated for IRF and SRF, but not PED nor SHRM, as these are not relevant for diabetic macular edema patients. More details on these datasets and the annotation protocol can be found in [27].
Labeled data exists for all 3 domains as detailed in Table S1. For each experimental setting, different ablation combinations were performed by training models on labeled source domain(s) $D^{s}$ and applying the trained models on unlabeled domain(s) $D^{t}$. When a domain is considered $D^{t}$, we omit its labels in the training and use them only for ($i$) evaluating the models on the test set of that domain, or ($ii$) training of an UpperBound model used as a reference. Our data stratification is detailed in Table S1.
@@ IV Results and Discussion
In most experimental settings, we train a model $F$ concurrently on $(x,y)\in D^{\mathrm{s}}$ and $x\in D^{\mathrm{t}}$, and then apply this trained model on the target data for evaluating $F(x\,|\,x\!\in\!D^{\mathrm{t}})\rightarrow y\in D^{\mathrm{t}}$. In the zero-shot DA setting, we train only on $(x,y)\in D^{\mathrm{s}}$, without any information from $D^{\mathrm{t}}$. We evaluate the trained models also on their initial source domain $F(x\,|\,x\!\in\!D^{\mathrm{s}})\rightarrow y\in D^{\mathrm{s}}$ to assess the retention of source-domain segmentation capability.

### v3

# Joint semi-supervised and contrastive learning enables domain generalization and multi-domain segmentation

@@ Abstract
Despite their effectiveness, current deep learning models face challenges with images coming from different domains with varying appearance and content. We introduce SegCLR, a versatile framework designed to segment images across different domains, employing supervised and contrastive learning simultaneously to effectively learn from both labeled and unlabeled data. We demonstrate the superior performance of SegCLR through a comprehensive evaluation involving three diverse clinical datasets of 3D retinal Optical Coherence Tomography (OCT) images, for the slice-wise segmentation of fluids with various network configurations and verification across 10 different network initializations. In an unsupervised domain adaptation context, SegCLR achieves results on par with a supervised upper-bound model trained on the intended target domain. Notably, we discover that the segmentation performance of SegCLR framework is marginally impacted by the abundance of unlabeled data from the target domain, thereby we also propose an effective domain generalization extension of SegCLR, known also as zero-shot domain adaptation, which eliminates the need for any target domain information. This shows that our proposed addition of contrastive loss in standard supervised training for segmentation leads to superior models, inherently more generalizable to both in- and out-of-domain test data. We additionally propose a pragmatic solution for SegCLR deployment in realistic scenarios with multiple domains containing labeled data. Accordingly, our framework pushes the boundaries of deep-learning based segmentation in multi-domain applications, regardless of data availability —– labeled, unlabeled, or nonexistent.
@@ I Introduction
Domain adaptation techniques aim to maintain a similar performance on images from new domains while reducing the annotation effort required in such newer datasets. Unsupervised domain adaptation aims to leverage information learned from a labeled data domain for applications in other domains where only unlabeled data is available. To this end, many deep learning methods have been proposed [8], mostly using generative adversarial networks, e.g., to translate visual appearance across OCT devices [9]. Ultimately, it is desirable for learned models to implicitly generalize to other domains without requiring any new data, i.e., ideally not even unlabeled data from a new domain. This scenario, known as domain generalization [10] or zero-shot domain adaptation [11], provides huge advantages in real-world deployment by obviating any retraining or customization being required per test domain encountered.
Advances in self-supervised learning have been increasingly successful in extracting informative features from images without accessing their labels, hence enabling the exploitation of much larger unlabeled datasets [12, 13, 14, 15, 16, 17, 18, 19]. These methods usually employ different versions of Siamese networks [20] that can be applied to different inputs in parallel while sharing their weights. For self-supervision, typically different augmentations of an image are treated as positive samples, which are desired to map close-by or to a single point on a learned manifold, and therefore minimizing the feature distances between these positive samples when passed through a Siamese network. To prevent such optimization from collapsing to a constant output, different approaches have been proposed: For instance, BYOL [16] and SimSiam [15] employ a stop-gradient operation for one of their two encoders, with their main difference being how the other encoder is updated: BYOL uses the moving average of the parameters of the stop-gradient encoder, whereas SimSiam uses a Siamese architecture between its two encoders. Another successful strategy, contrastive learning (CL), introduces the concept of negative pairs, the distances of which should meanwhile be maximized. SimCLR [13] is one of the most widely-adopted CL frameworks. Self-supervised learning is commonly used for model pretraining, typically based on natural images such as ImageNet [21]. Such pretrained model is subsequently fine-tuned or distilled for downstream tasks such as classification, detection, or segmentation [14]. Notably, while BYOL is a more recent approach, when handling multiple domains, better results have been reported [22] with SimCLR, on which we build our method in this work.
@@
We propose SegCLR as a CL solution to improve segmentation quality by leveraging both manual annotations and unlabeled data. To that end we introduce a semi-supervised framework for joint training of CL together with segmentation labels. Preliminary results of this work have been presented as a conference contribution [25] for slice-wise segmentation of two datasets acquired with different OCT scanners, where the domain shift originates from a change in acquisition device. SegCLR without any target-domain labels is shown to achieve unsupervised domain adaptation results close to an upper bound that uses supervised learning on the target domain. Our additional contributions in this work are: (1) In addition to acquisition device change, we present results with a third dataset with a new disease indication, which helps demonstrate the benefits of SegCLR also for domain shift in disease type. (2) An analysis of the amount of unlabeled target data is presented, indicating its limited effect on the resulting segmentation performance. (3) Exploiting the finding above, we propose a domain generalization use-case of our SegCLR framework, for which an extensive analysis reveals substantial benefits showing that SegCLR trained models are implicitly better at generalizing to unseen domains. (4) An analysis of the stability of the results upon distinct network initialisations reveals that most proposed SegCLR variants perform well and similarly, confirming the stability of the overall SegCLR framework to method setting choices such as pair generation and contrastive projection strategies. This further highlights the importance of analyzing many randomized result replicates in assessing deep learning models in general. (5) In a multi-domain training scenario of SegCLR, we show that with labeled data available from multiple domains, SegCLR applied to the domains simultaneously yields results superior to conventional supervised training as well as SegCLR trained on individual domains, potentially better leveraging cross-domain information.
@@ II Methods
The considered learning frameworks are illustrated in Figure 1. Let a source domain ($D^{s}$) contain labeled data, with which Baseline models can be trained following typical supervised learning schemes. Distinct domains for which only unlabeled images can be observed are considered target domains ($D^{t}$). We first study an unsupervised domain adaptation setting as in [25], where SegCLR is trained for a source domain $D^{s}$ and a target domain $D^{t}$ by using a combination of supervision and contrastive losses. We then investigate its domain generalization adaptation application, where SegCLR is trained using only data from a labeled source domain $D^{s}$, while being applied on unseen target domains $D^{t}$, i.e., on domains for which no data (labeled or unlabeled) is used in training.
Figure 1: Illustration of Unsupervised Domain Adaptation (UDA) and Domain Generalization frameworks studied herein for SegCLR. The colored ellipses indicate the losses to use in training, with the variable superscripts representing the domain being source (s) or target (t).
@@ II-A Supervised learning
For the segmentation backbone, we adopted the proven UNet architecture [2] (details in Figure  2), which can be modeled as $F(\cdot)$ processing an image $x$ to predict a segmentation map $p=F(x)$ to approximate an (expert-annotated) ground truth segmentation $y$. Each segmentation map contains $C$ classes, for which $y^{c}$ and $p^{c}$ represent, respectively, the ground truth and predicted probabilities of class $c$. The model $F$ learns to segment all classes present in the image by minimizing a supervised loss $\mathcal{L}_{\mathrm{sup}}$, which in our work is the logarithmic Dice loss of labeled data in a source domain $D^{s}$, i.e.,
$\mathcal{L}_{\mathrm{sup}}=-\frac{1}{C}\sum_{c=1}^{C}\sum_{(p_{i},y_{i})\in D^{s}}\log{\frac{2\sum_{j\in\mathrm{pixels}}y_{i}^{c}p_{i}^{c}}{\epsilon+\sum_{j\in\mathrm{pixels}}(y_{i}^{c}+p_{i}^{c})}}$ (1)
for all training images $(x_{i},y_{i})\in D^{s}$, where $\epsilon$ is a small number ($10^{-12}$) to avoid division by zero. Note that volumetric OCT data is processed slice-wise for 2D segmentation.
@@ II-B Self-supervised learning
Self-supervised learning aims to optimize an encoder $E(\cdot)$ to achieve good representations $h=E(x)$ without the need of manually annotated labels $y$. In the literature (for mostly classification tasks), the ResNet architecture is commonly used for $E(\cdot)$. Instead, we use the existing UNet encoder in order to adapt the learned features $h$ for our intended segmentation task. A subsequent contrastive projection head $C(\cdot)$ then maps the bottleneck-layer features to vector projections $z=C(h)$ on which the contrastive loss $\mathcal{L}_{\mathrm{con}}$ is applied. The architecture used for $E(\cdot)$ and $C(\cdot)$ is illustrated in Figure 2. We herein employ two widely adopted self-supervised learning frameworks: SimCLR [13] and SimSiam [15].
@@ II-C Joint semi-supervised and contrastive learning
We adapt the USCL joint training strategy, which was proposed for US video classification, to our segmentation task by combining $\mathcal{L}_{\mathrm{sup}}$ and $\mathcal{L}_{\mathrm{con}}$ in a semi-supervised framework illustrated in Figure 3. Considering a source domain $D^{\mathrm{s}}$ and a target domain $D^{\mathrm{t}}$, total loss $\mathcal{L}$ is calculated as follows:
@@
Figure 3: Illustration of semi-supervised contrastive learning framework for unsupervised domain adaptation. The SegCLR block corresponds to the architecture in Figure 2. The repel losses are not used by SimSiam. Supervised losses are only used when labeled images exist. While this framework is flexible to accommodate any number of labeled images, at least one is required to drive the decoder arm of the underlying UNet.
@@ II-D Pair generation strategies
The approach chosen for contrastive pair generation is key in self-supervised learning. We herein propose and compare different pair generation functions $P(\cdot)$ for volumetric OCT images, as illustrated in Figure 4.
First, the augmentation-based pair formation typically employed for natural images (e.g., in SimCLR and SimSiam) is adapted for OCT slices, denoted as $P_{\mathrm{a}}$. To that end, labeled slices in $D^{\mathrm{s}}$ and random slices in $D^{\mathrm{t}}$ are augmented with horizontal flipping ($p=0.5$), horizontal and vertical translation (within 25% of the image size), zoom in (up to 50%), and color distortion (brightness up to 60% and jittering up to 20%). For color augmentation, images are first transformed to RGB and then back to grayscale.
@@ II-F Implementation
Adam optimizer [26] was used in all models, with a learning rate of $10^{-3}$. Dropout with $p=0.5$ is applied on the layers illustrated with orange arrows in Figure 2. The projection head $\rho^{\textrm{MLP}}$ in $C(\cdot)$ is formed by two fully-connected layer with 128 units each, where the first one uses group normalization [27] (with a group size of 4) and ReLU activation. We heuristically set $\lambda=20$ and the standard deviation for $P_{\mathrm{s}}$ as $\sigma=250\,\mu m$, which is the range for which we observe roughly similar features across slices. Supervised models were trained for 200 epochs, and the model at the epoch with the highest average Dice coefficient across classes on the validation set was selected for evaluation on a holdout test set. Our implementation is in Tensorflow 2.7, ran on an NVIDIA V100 GPU.
@@ II-G Evaluation setup
Individual slices of the 3D OCT volumes are segmented with UNet in 2D. All reported experiments were replicated by training from scratch with 10 different initialization seeds, to factor out the sensitivity to randomness in network initialization. Model performance was then also evaluated slice-wise and class-wise using slices with ground-truth annotations. The Dice coefficient serves as an overlap metric. In the situation that the ground truth and prediction maps are both empty, a perfect Dice score is given to that class on that slice. The Unnormalized Volume Dissimilarity (UVD) is also included as an error metric for the total segmentation error (FP+FN), i.e., the sum of any (mis)segmented volume that is outside the ground truth and any ground truth volume that was not segmented by a method. Compared to Dice, UVD is more robust to FP on B-scans with small annotated regions for individual classes. Dice is reported as % (higher is better) and UVD as femtoliters (fL) (lower is better).
Visualizing results by averaging metrics across classes with a large variation may lead to bias. Thus, for the metrics reported as boxplots, each per-slice metric $m^{c}_{i}$ for method $i$ and class $c$ is first normalized by its class Baseline $m^{c}_{\textrm{bas}}$, and then these are averaged over all $c$ and images on the test set. The average of all Baseline replicates is used as $m^{c}_{\textrm{bas}}$ for calculating the relative metrics of all other models. Absolute metrics for each class are also reported in the tables as mean $\pm$ standard deviation. To facilitate model comparison, the metrics included in most tables in the main text are averaged across classes. To further allow for the interpretation and comparison of absolute segmentation performances, we included detailed tables in Supplementary Materials with metrics separately for individual labels.
Statistical significance is assessed using a paired t-test and reported in the boxplots with **** for $\leq$0.0001, *** for $\leq$0.001, ** for $\leq$0.01, * for $\leq$0.05 (*), and ‘n.s.’ for non-significant. When no bar joins two result boxes, the marker indicates significance for the test performed against the corresponding baseline.
For comparing and ranking models, we adapted the strategy from [28] to simultaneously consider both Dice and UVD as complementary metrics. Accordingly, for each individual volumetric eye image, the models are first ranked among themselves for each initialization seed, separately based on their Dice and UVD. These rankings are next averaged across the 10 seed replicates of all the volumetric images as well as across the two metrics, to obtain and report a final ranking of each model for a given test domain.
@@ III Dataset
Selected B-scans from $D_{1}$ and $D_{2}$ were manually annotated for fluid regions of potential diagnostic value, i.e., intraretinal fluid (IRF), subretinal fluid (SRF), pigment epithelial detachment (PED), and subretinal hyperreflective material (SHRM). B-scans from $D_{3}$ were annotated for IRF and SRF, but not PED nor SHRM, as these are not relevant for diabetic macular edema patients. More details on these datasets and the annotation protocol can be found in [29].
For computations in this work, all slices (B-scans) from the three datasets were resampled to $512\times 512$ pixels using linear interpolation, yielding the approximate resolution of $10\times 4$ $\mu$m/pixel. Corresponding ground-truth labels were also resampled accordingly using nearest neighbours.
Labeled data exists for all 3 domains as detailed in Table S1 together with the stratification details. For each experimental setting as described in the corresponding following sections, different ablation combinations were performed by training models on labeled source domain(s) $D^{s}$ and applying (testing) such trained models on unlabeled target domain(s) $D^{t}$. When a domain is considered $D^{t}$, we omit its labels in the training and use them only for ($i$) evaluating the models on the test set of that domain, or ($ii$) training of an UpperBound model used as a reference. The train/test configuration of each subsequent section is also summarized in Table S2.
@@ IV Results and Discussion
In most experimental settings, we train a model $F$ concurrently on $(x,y)\in D^{\mathrm{s}}$ and $x\in D^{\mathrm{t}}$, and then apply this trained model on the target data for evaluating $F(x\,|\,x\!\in\!D^{\mathrm{t}})\rightarrow y\in D^{\mathrm{t}}$. In the domain generalization setting, we train only on $(x,y)\in D^{\mathrm{s}}$, without any information from $D^{\mathrm{t}}$. We evaluate the trained models also on their initial source domain $F(x\,|\,x\!\in\!D^{\mathrm{s}})\rightarrow y\in D^{\mathrm{s}}$ to assess the retention of source-domain segmentation capability.